# Demagnetization in micromagnetics: magnetostatic self-interactions of bulk chiral magnetic skyrmions

### v1

@@ Abstract
It is known that the magnetization in a ferromagnet induces an internal demagnetizing magnetic field. This, in-turn, generates a magnetostatic self-energy, coming from dipole-dipole interactions. In general, this energy is singular and intrinsically non-local. However, one can introduce a magnetic scalar potential associated with the demagnetizing field. We show that computation of the magnetic potential can be achieved by solving Poisson’s equation where the source is the divergence of the magnetization. Furthermore, we determine the variation of the magnetostatic self-energy and its associated back-reaction on the magnetization. Then, we present a theoretical study of the effect of the long-range dipolar interaction on magnetic skyrmion textures in the bulk of chiral ferromagnetic materials with the Dzyaloshinskii–Moriya interaction. It is observed that the demagnetizing dipolar self-interaction can stabilize antiskyrmion crystals in Heusler compounds.
@@ I Introduction
Topological spin textures arise in magnetic systems due to the competition between the Heisenberg exchange interaction and the Dzyaloshinskii–Moriya interaction (DMI). The Heisenberg exchange energy promotes parallel alignment of spins whereas the DMI energy favours non-collinear alignment of spins [1]. Magnetic skyrmions are one such type of topological spin texture. They are topological defects in the magnetization configuration and behave as magnetic quasi-particles. Such vortex-like configurations were first observed in [2, 3] in the context of condensed matter systems and, in particular, they found skyrmions in easy-axis chiral magnetic materials with the Dzyaloshinskii–Moriya interaction. The easy-axis, or Zeeman, interaction is essential for stability as it provides a scale for the magnetic skyrmion.
@@
There are three different skyrmion types investigated in the literature, each associated to a particular DMI term. The first two are the Bloch and Néel skyrmions, corresponding to the Dresselhaus and Rashba spin-orbit couplings (SOC), respectively. The final type is the antiskyrmion, which appears in Heusler compounds [4, 5, 6], so we label the associated DMI term as the Heusler DMI. The chiral magnet can also be formulated as a $\textup{U}(1)\times\textup{SU}(2)$ gauge theory with a constant background gauge field [7]. Within this formulation, the three DMI terms are related by gauge transformations, which are controlled by a single rotation parameter [8]. The energy functional is invariant under the relevant gauge transformations, so the three skyrmion types are degenerate in energy. An in-depth review on the current state of the theoretical and experimental realization of magnetic skyrmions in condensed matter systems can be found in [9].
@@
The inclusion of the magnetic dipole-dipole interaction (DDI) introduces frustration into a ferromagnetic state, turning it into a frustrated spin system. This frustration between the DDI and Heisenberg exchange interaction can lead to interesting phenomena. The nucleation of skyrmions using an external magnetic dipole, which modifies the Zeeman interaction term, was demonstrated numerically by Garanin et al. [10]. Depending on how the magnetic dipole approaches the film, a skyrmion-antiskyrmion pair can be created in the bulk or a skyrmion can be created at the edge of the film. It was also shown that magnetic biskyrmions can be stabilized by dipolar interactions in centrosymmetric materials where the DMI term is forbidden [11].
@@
Kashuba and Pokrovsky [12] presented a theory on the effect of long-range dipole interactions in thin ferromagnetic films. In particular, they studied stripe domain structures in ordinary ferromagnets with no interfacial DMI. An analytical investigation into the stabilization of magnetic skyrmions by dipolar interactions in thin films was conducted by Ezawa [13]. He considered both ordinary and chiral ferromagnets, and employed the Dresselhaus DMI term in the latter, which favours Bloch skyrmions. In both ferromagnet types, domain wall stripes and skyrmions were constructed analytically. However, in his analysis, Ezawa considers the spin texture of a Néel skyrmion, which is incompatible with the Dresselhaus DMI.
@@
A numerical study on the effect of the dipolar interaction on chiral ferromagnets was performed by Kwon at al. [14]. They also considered the Dresselhaus DMI term, for which they investigated stripe domains and Bloch skyrmion lattices in the presence of an applied external magnetic field and perpendicular uniaxial anisotropy. As the external magnetic field is increased, the microscopic spin configurations transform from a spin-spiral stripe to a skyrmion crystalline structure. Sui and Hu [15] conducted a similar investigation (but with no anisotropies), wherein they considered the DMI term corresponding to Rashba SOC, which gives rise to Néel type skyrmions. In both studies [14, 15], magnetic skyrmion crystals were investigated using simulated annealing with a Monte-Carlo method on small grids. Furthermore, they both include a tuneable constant that allows them to alter the strength of the dipolar interaction, and Kwon et al. chose to set the dipolar interaction strength comparable to the anisotropy. Sui and Hu [15] find that, as the DDI strength increases, the skyrmion crystal structure changes from a hexagonal arrangement to one of a square form.
@@
Our approach differs from that of [14, 15]. Instead of considering a tuneable dipole interaction parameter, we focus on the effect the dipolar interaction has on the different DMI-related skyrmion types in the bulk of a chiral ferromagnet. We consider three DMI terms and their associated skyrmion type - Dresselhaus (Bloch), Rashba (Néel) and Heusler (antiskyrmion). Without the dipolar interaction, the magnetic skyrmion types are equivalent up to redefinitions of the field. In addition to this, the preferred crystalline configuration is hexagonal for all three DMI terms. However, the dipole-dipole interaction breaks this equivalence and there is a splitting in the energies. Furthermore, the axial symmetry is broken for the antiskyrmion (Heusler DMI), resulting in a change in shape of both isolated antiskyrmions, and their crystal structure.
@@ II The model
$N_{\textup{Sk}}=\frac{1}{4\pi}\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,\mathbf{n}\cdot\left(\frac{\partial\mathbf{n}}{\partial x_{1}}\times\frac{\partial\mathbf{n}}{\partial x_{2}}\right).$ (5)
Following the usual convention of the literature, magnetic skyrmions with $N_{\textup{Sk}}<0$ are referred to as skyrmions and antiskyrmions for $N_{\textup{Sk}}>0$. The skyrmion type is determined by the choice of DMI term. Magnetic skyrmions with $N_{\textup{Sk}}=0$, such as a chiral magnetic bag enclosing a domain wall, are referred to as skyrmionium [16] or a $2\pi$-vortex [17]. We will, however, not consider skyrmion bags in this article.
@@
where we have introduced the Lifshitz invariants [18],
@@
The final DMI term we consider arises from the vectors $\{\mathbf{d}_{1}=\mathbf{e}_{1},\,\mathbf{d}_{2}=-\mathbf{e}_{2}\}$ and corresponds to tetragonal compounds with $D_{2d}$ symmetry (such as Heusler compounds) [19],
@@
where $f(r)$ decreases monotonically from $\pi$ to 00. Note that this field has degree $N_{\textup{Sk}}=-1$, so we call it a skyrmion, following the standard convention in the magnetism literature. It follows from (10) that the Dresselhaus DMI term favours Bloch skyrmions
@@
The magnetization vector $\mathbf{n}\in S^{2}$ can be represented graphically with a HSV colour scheme associated to the Runge colour sphere, and is presented as laid out in [20]. The phase $\arg(n_{1}+in_{2})$ determines the hue where $\arg(n_{1}+in_{2})=0$ is identified with red, $\arg(n_{1}+in_{2})=2\pi/3$ with green and $\arg(n_{1}+in_{2})=4\pi/3$ with blue [21]. Finally, the value of $n_{3}$ determines the brightness [22], such that spin-up states $\mathbf{n}_{\uparrow}=(0,0,1)$ are white and the spin-down states $\mathbf{n}_{\downarrow}=(0,0,-1)$ are black.
@@
$\bm{\nabla}\cdot\mathbf{n}$ 00 $f^{\prime}(r)\cos f(r)+r^{-1}\sin f(r)$ $\left(-f^{\prime}(r)\cos f(r)+r^{-1}\sin f(r)\right)\sin 2\theta$
Table 1: Summary of the three different DMI terms we consider and their corresponding skyrmion type. Also shown is the divergence of each skyrmion, which is relevant as it is the source $\rho=-(\bm{\nabla}\cdot\mathbf{n})$ in the Poisson equation $\Delta\psi=\mu\rho$ for the magnetic potential.
@@ III The interaction energy of a distribution of magnetic dipoles
In this section we define the magnetostatic self-energy, or dipole-dipole interaction energy, $E_{\textup{DDI}}$ and compute its first variation. The magnetic field induced by an isolated dipole of moment $\mathbf{m}$ at $\mathbf{0}$ is [23]
@@
Comparing with (18), we see that the magnetic potential $\psi$ satisfies the PDE [24]
@@
that is, Poisson’s equation with source $\rho=-(\bm{\nabla}\cdot\mathbf{m})$. Up to a multiplicative constant, the magnetic field induced by the dipole distribution $\mathbf{m}$ coincides, therefore, with the magnetic field induced by the charge distribution $-(\bm{\nabla}\cdot\mathbf{m})$ [25]. Roughly speaking, we may think of $-(\bm{\nabla}\cdot\mathbf{m})$ as “magnetic charge density”.
The interaction energy of a pair of magnetic dipoles $\mathbf{m}^{(1)}$, $\mathbf{m}^{(2)}$, is $-\mathbf{m}^{(1)}\cdot\mathbf{B}^{(2)}$, where $\mathbf{B}^{(2)}$ is the magnetic field induced by $\mathbf{m}^{(2)}$ at the position of $\mathbf{m}^{(1)}$. Hence, the total dipole-dipole interaction energy of a continuous dipole density distribution is the magnetostatic energy
@@
=\,= $-\frac{1}{2}\int_{\Omega}\textup{d}^{3}\mathbf{x}\,\left(\bm{\nabla}\cdot\mathbf{m}\right)\psi+\frac{1}{2}\int_{\partial\Omega}\textup{d}\mathbf{s}\cdot\left(\psi\mathbf{m}\right)$
=\,= $\frac{1}{2\mu_{0}}\int_{\Omega}\textup{d}^{3}\mathbf{x}\,\psi\Delta\psi+\frac{1}{2}\int_{\partial\Omega}\textup{d}\mathbf{s}\cdot\left(\psi\mathbf{m}\right),$ (23)
by the Divergence Theorem. We will use this formula only in situations where the boundary conditions ensure that the boundary term vanishes. In this case, $E_{\textup{DDI}}$ coincides with the “electrostatic” self-energy of the “charge” distribution $-(\bm{\nabla}\cdot\mathbf{m})$.
[…]
@@ IV Evading the Hobart–Derrick theorem
So far we have shown how to include the dipolar interaction and computed its variation with respect to the magnetization. Now, we want to understand how the dipolar interaction transforms under a coordinate rescaling, to see the role it plays in evading the Hobart–Derrick theorem [26].
@@
With the dimensionless total energy (31) in hand, we now turn to evading the Hobart–Derrick theorem [26]. Let us consider, for a given field $\mathbf{n}:{\mathbb{R}}^{2}\rightarrow S^{2}$, the one parameter variation $\mathbf{n}_{\lambda}(\mathbf{x}):=\mathbf{n}(\lambda\mathbf{x})$, $\lambda>0$, obtained by isotropic dilation. The “magnetic charge density” of $\mathbf{n}_{\lambda}$ is $\rho_{\lambda}(x)=-\bm{\nabla}\cdot\mathbf{n}_{\lambda}(\mathbf{x})=-\lambda\rho(\lambda\mathbf{x})$, so its magnetic potential $\psi_{\lambda}$ satisfies

### v2

@@ Abstract
We develop a theoretical and numerical framework for three-dimensional bulk chiral magnets that includes the full magnetostatic dipole-dipole interaction and its back-reaction on the magnetization. Assuming translational invariance along one spatial direction, we analyze the effect of dipolar interactions on three Dzyaloshinskii–Moriya interaction (DMI) terms – Dresselhaus, Rashba, and Heusler –corresponding to Bloch, Néel, and antiskyrmion textures. In the absence of the dipolar interaction, these three DMI terms are gauge-equivalent and yield degenerate skyrmion energies. Incorporating the non-local dipole-dipole interaction breaks this degeneracy: Bloch skyrmions remain unaffected, Néel skyrmions shrink slightly, and Heusler antiskyrmions lose axial symmetry and stabilize into square-lattice crystals. The system is solved using a non-local numerical relaxation method that self-consistently computes the magnetostatic potential from Poisson’s equation. Our results show that long-range dipolar interactions can stabilize bulk antiskyrmion crystals in translationally invariant three-dimensional chiral magnets.
@@ I Introduction
Topological spin textures arise in magnetic systems due to the competition between the Heisenberg exchange interaction and the Dzyaloshinskii–Moriya interaction (DMI). The Heisenberg exchange energy promotes parallel alignment of spins whereas the DMI energy favours non-collinear alignment of spins [Muratov_2020]. Magnetic skyrmions are one such type of topological spin texture. They are topological defects in the magnetization configuration and behave as magnetic quasi-particles. Such vortex-like configurations were first observed in [Bogdanov_1989, Bogdanov_1994] in the context of condensed matter systems and, in particular, they found skyrmions in easy-axis chiral magnetic materials with the Dzyaloshinskii–Moriya interaction. The easy-axis, or Zeeman, interaction is essential for stability as it provides a scale for the magnetic skyrmion.
@@
There are three different skyrmion types investigated in the literature, each associated to a particular DMI term. The first two are the Bloch and Néel skyrmions, corresponding to the Dresselhaus and Rashba spin-orbit couplings (SOC), respectively. The final type is the antiskyrmion, which appears in Heusler compounds [Saha_2019, Felser_2022, Peng_2020], so we label the associated DMI term as the Heusler DMI. The chiral magnet can also be formulated as a $\textup{U}(1)\times\textup{SU}(2)$ gauge theory with a constant background gauge field [Amari_Nitta_2023]. Within this formulation, the three DMI terms are related by gauge transformations, which are controlled by a single rotation parameter [Schroers_2020]. The energy functional is invariant under the relevant gauge transformations, so the three skyrmion types are degenerate in energy. An in-depth review on the current state of the theoretical and experimental realization of magnetic skyrmions in condensed matter systems can be found in [Back_2020].
@@
The inclusion of the magnetic dipole-dipole interaction (DDI) introduces frustration into a ferromagnetic state, turning it into a frustrated spin system. This frustration between the DDI and Heisenberg exchange interaction can lead to interesting phenomena. The nucleation of skyrmions using an external magnetic dipole, which modifies the Zeeman interaction term, was demonstrated numerically by Garanin et al. [Garanin_2018]. Depending on how the magnetic dipole approaches the film, a skyrmion-antiskyrmion pair can be created in the bulk or a skyrmion can be created at the edge of the film. It was also shown that magnetic biskyrmions can be stabilized by dipolar interactions in centrosymmetric materials where the DMI term is forbidden [Gobel_2019].
@@
Kashuba and Pokrovsky [Kashuba_1993] presented a theory on the effect of long-range dipole interactions in thin ferromagnetic films. In particular, they studied stripe domain structures in ordinary ferromagnets with no interfacial DMI. An analytical investigation into the stabilization of magnetic skyrmions by dipolar interactions in thin films was conducted by Ezawa [Ezawa_2010]. He considered both ordinary and chiral ferromagnets, and employed the Dresselhaus DMI term in the latter, which favours Bloch skyrmions. In both ferromagnet types, domain wall stripes and skyrmions were constructed analytically. However, in his analysis, Ezawa considers the spin texture of a Néel skyrmion, which is incompatible with the Dresselhaus DMI.
@@
A numerical study on the effect of the dipolar interaction on chiral ferromagnets was performed by Kwon at al. [Kwon_2012]. They also considered the Dresselhaus DMI term, for which they investigated stripe domains and Bloch skyrmion lattices in the presence of an applied external magnetic field and perpendicular uniaxial anisotropy. As the external magnetic field is increased, the microscopic spin configurations transform from a spin-spiral stripe to a skyrmion crystalline structure. Sui and Hu [Sui_2022] conducted a similar investigation (but with no anisotropies), wherein they considered the DMI term corresponding to Rashba SOC, which gives rise to Néel type skyrmions. In both studies [Kwon_2012, Sui_2022], magnetic skyrmion crystals were investigated using simulated annealing with a Monte-Carlo method on small grids. Furthermore, they both include a tuneable constant that allows them to alter the strength of the dipolar interaction, and Kwon et al. chose to set the dipolar interaction strength comparable to the anisotropy. Sui and Hu [Sui_2022] find that, as the DDI strength increases, the skyrmion crystal structure changes from a hexagonal arrangement to one of a square form.
@@
Our approach differs from that of [Kwon_2012, Sui_2022]. Instead of considering a tuneable dipole interaction parameter, we focus on the effect the dipolar interaction has on the different DMI-related skyrmion types in the bulk of a chiral ferromagnet. That is, we do not model a thin film, instead we model a three dimensional bulk system and impose translation invariance along the $z$-direction. We consider three DMI terms and their associated skyrmion type - Dresselhaus (Bloch), Rashba (Néel) and Heusler (antiskyrmion). Without the dipolar interaction, the magnetic skyrmion types are equivalent up to redefinitions of the field. However, the dipole-dipole interaction breaks this equivalence and there is a splitting in the energies. Furthermore, the axial symmetry is broken for the antiskyrmion (Heusler DMI), resulting in a change in shape of both isolated antiskyrmions, and their crystal structure.
@@ II The model
$N_{\textup{Sk}}=\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,\rho(\mathbf{x}),\quad\rho(\mathbf{x})=\frac{1}{4\pi}\mathbf{n}\cdot\left(\frac{\partial\mathbf{n}}{\partial x_{1}}\times\frac{\partial\mathbf{n}}{\partial x_{2}}\right).$ (5)
Following the usual convention of the literature, magnetization fields with $N_{\textup{Sk}}<0$ are referred to as skyrmions while those with for $N_{\textup{Sk}}>0$ are antiskyrmions. The skyrmion type is determined by the choice of DMI term. Magnetic skyrmions with $N_{\textup{Sk}}=0$, such as a chiral magnetic bag enclosing a domain wall, are referred to as skyrmionium [Rybakov_2020] or a $2\pi$-vortex [Bogdanov_1999]. We will, however, not consider skyrmion bags in this article.
In addition to magnetic skyrmions having a topological degree, they also have an associated size [Wang_2018]. There is some choice in how this is defined [Adam_2016]. For a field theory permitting topological solitons, one can introduce the root-mean-square (RMS) radius of the soliton, defined by
$R_{\textup{Sk}}=\left(\frac{1}{N_{\textup{Sk}}}\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,r^{2}|\rho(\mathbf{x})|\right)^{\frac{1}{2}}.$ (6)
We will use this definition of skyrmion size throughout to compare what happens to the size of the skyrmion as the DDI is included for various DMI terms.
@@
where we have introduced the Lifshitz invariants [Leonov_2016],
@@
The final DMI term we consider arises from the vectors $\{\mathbf{d}_{1}=\mathbf{e}_{1},\,\mathbf{d}_{2}=-\mathbf{e}_{2}\}$ and corresponds to tetragonal compounds with $D_{2d}$ symmetry (such as Heusler compounds) [Nayak_2017],
@@
where $f(r)$ decreases monotonically from $\pi$ to 0. Note that this field has degree $N_{\textup{Sk}}=-1$, so we call it a skyrmion, following the standard convention in the magnetism literature. It follows from (11) that the Dresselhaus DMI term favours Bloch skyrmions
@@
The magnetization vector $\mathbf{n}\in S^{2}$ can be represented graphically with a HSV colour scheme associated to the Runge colour sphere, and is presented as laid out in [Leask_2022]. The phase $\arg(n_{1}+in_{2})$ determines the hue where $\arg(n_{1}+in_{2})=0$ is identified with red, $\arg(n_{1}+in_{2})=2\pi/3$ with green and $\arg(n_{1}+in_{2})=4\pi/3$ with blue [Barsanti_2020]. Finally, the value of $n_{3}$ determines the brightness [Rybakov_2019], such that spin-up states $\mathbf{n}_{\uparrow}=(0,0,1)$ are white and the spin-down states $\mathbf{n}_{\downarrow}=(0,0,-1)$ are black.
@@
$\bm{\nabla}\cdot\mathbf{n}$ 0 $f^{\prime}(r)\cos f(r)+r^{-1}\sin f(r)$ $\left(-f^{\prime}(r)\cos f(r)+r^{-1}\sin f(r)\right)\sin 2\theta$
Table 1: Summary of the three different DMI terms we consider and their corresponding skyrmion type. Here, we consider only axially symmetric configurations without the dipole-dipole interaction (DDI) present. Also shown is the divergence of each skyrmion $\bm{\nabla}\cdot\mathbf{n}$, which is relevant as it is the source $\rho=-(\bm{\nabla}\cdot\mathbf{n})$ in the Poisson equation $\Delta\psi=\mu\rho$ for the magnetic potential, and it provides information about the nature of DDI. For example, we expect the Bloch skyrmion to be unaffected by the DDI, the magnetic potential for the Néel skyrmion to be axially symmetric and the Heusler antiskyrmion to lose be axial symmetry.
@@ III The interaction energy of a distribution of magnetic dipoles
In this section we define the magnetostatic self-energy, or dipole-dipole interaction energy, $E_{\textup{DDI}}$ and compute its first variation. The magnetic field induced by an isolated dipole of moment $\mathbf{m}$ at $\mathbf{0}$ is [Qin_2018]
@@
Comparing with (19), we see that the magnetic potential $\psi$ satisfies the PDE [Fratta_2020]
@@
that is, Poisson’s equation with source $\rho=-(\bm{\nabla}\cdot\mathbf{m})$. Up to a multiplicative constant, the magnetic field induced by the dipole distribution $\mathbf{m}$ coincides, therefore, with the magnetic field induced by the charge distribution $(\bm{\nabla}\cdot\mathbf{m})$ [Mozurkewich_1979]. Roughly speaking, we may think of $(\bm{\nabla}\cdot\mathbf{m})$ as “magnetic charge density”.
The interaction energy of a pair of magnetic dipoles $\mathbf{m}^{(1)}$, $\mathbf{m}^{(2)}$, is $\mathbf{m}^{(1)}\cdot\mathbf{B}^{(2)}$, where $\mathbf{B}^{(2)}$ is the magnetic field induced by $\mathbf{m}^{(2)}$ at the position of $\mathbf{m}^{(1)}$. Hence, the total dipole-dipole interaction energy of a continuous dipole density distribution is the magnetostatic energy
@@
$=\,$ $-\frac{1}{2}\int_{\Omega}\textup{d}^{3}\mathbf{x}\,\left(\bm{\nabla}\cdot\mathbf{m}\right)\psi+\frac{1}{2}\oint_{\partial\Omega}\textup{d}\mathbf{s}\cdot\left(\psi\mathbf{m}\right)$
$=\,$ $\frac{1}{2\mu_{0}}\int_{\Omega}\textup{d}^{3}\mathbf{x}\,\psi\Delta\psi+\frac{1}{2}\oint_{\partial\Omega}\textup{d}\mathbf{s}\cdot\left(\psi\mathbf{m}\right),$ (24)
by the Divergence Theorem. We will use this formula only in situations where the boundary conditions ensure that the boundary term vanishes. In this case, $E_{\textup{DDI}}$ coincides with the “electrostatic” self-energy of the “charge” distribution $(\bm{\nabla}\cdot\mathbf{m})$.
We now specialize to the case of immediate interest: the dipole-dipole interaction energy of a chiral ferromagnet in a translation invariant configuration. The dipole field has constant length $M_{s}$, so $\mathbf{m}=M_{s}\mathbf{n}$ where $\mathbf{n}$ is valued on the unit sphere. Further, we impose translation invariance in the direction $\mathbf{e}_{3}=(0,0,1)$, so $\mathbf{n}$ is independent of $x_{3}$. The boundary conditions require some care. We assume that some anisotropy in the system (either intrinsic or generated by an external applied magnetic field) imposes an energetic preference for the dipole orientation $\mathbf{n}=\mathbf{e}_{3}$, and consider fields $\mathbf{n}:\mathbb{R}^{2}\rightarrow S^{2}$ which have compact support in the sense that there exists $R_{0}>0$ such that, for all $r:=|(x_{1},x_{2})|\geq R_{0}$, $\mathbf{n}(x_{1},x_{2})=\mathbf{e}_{3}$. Since the field $\mathbf{n}$ is translation invariant, the total dipole interaction energy either vanishes (for example, if $\mathbf{n}$ is constant) or diverges. The energy per unit length (in the $\mathbf{e}_{3}$ direction) may be finite however, and this coincides with the total energy of the slab $\Omega=\mathbb{R}^{2}\times[0,1]$, which we compute as the limit of the energy of the thick disk $\Omega_{R}=\{\mathbf{x}\>:\>x_{1}^{2}+x_{2}^{2}\leq R^{2},\>0\leq x_{3}\leq 1\}$ as $R\rightarrow\infty$. Note that, in this case, the boundary term in (III) vanishes identically for all $R>R_{0}$: the flux of $\psi\mathbf{m}$ through the cylindrical wall vanishes since $\mathbf{m}$ has no normal component on this part of the boundary, and the flux through the top disk (at $x_{3}=1$) is exactly canceled by the flux through the bottom disk (at $x_{3}=0$) since $\psi\mathbf{m}$ is translation invariant. Hence
@@ IV Evading the Hobart–Derrick theorem
So far we have shown how to include the dipolar interaction and computed its variation with respect to the magnetization. Now, we want to understand how the dipolar interaction transforms under a coordinate rescaling, to see the role it plays in evading the Hobart–Derrick theorem [Derrick_1964].
@@
With the dimensionless total energy (32) in hand, we now turn to evading the Hobart–Derrick theorem [Derrick_1964]. Let us consider, for a given field $\mathbf{n}:{\mathbb{R}}^{2}\rightarrow S^{2}$, the one parameter variation $\mathbf{n}_{\lambda}(\mathbf{x}):=\mathbf{n}(\lambda\mathbf{x})$, $\lambda>0$, obtained by isotropic dilation. The “magnetic charge density” of $\mathbf{n}_{\lambda}$ is $\rho_{\lambda}(x)=-\bm{\nabla}\cdot\mathbf{n}_{\lambda}(\mathbf{x})=-\lambda\rho(\lambda\mathbf{x})$, so its magnetic potential $\psi_{\lambda}$ satisfies